M W427L 93390 Summer’06

TWTh 11:30-1:00 ECJ 1.204; MF 11:30-1:00 ECJ 1.204 (93810; TA: Michael Ortiz [mortiz@math.utexas.edu]), MF 11:30-1:00 ETC 2.136 (93815; TA: Ricardo Teixeira [rteixeira@math.utexas.edu]),

*** Usually T W Th classes will be lectures, M F discussion sections.

Instructor: C. Friedman RLM 11.140 471-5161 office hrs: T W Th 2:00-3:00pm or by appt. Web Page (homework assignments are listed): www.ma.utexas.edu/˜ friedman TA: []

Text and Syllabus: Vector Calculus , 5th ed. Marsden and Tromba, Freeman Publ. Syllabus; Parts of Chapters I-VIII.( Various technical sections omitted or discussed only briefly. Exactly which sections will be covered, depends on how quickly we progress, etc.)

Exams: We will have exams on 6/27 (class 11; Tues), 7/19 (class 20; Wed), 8/9 (Class 29; Wed); no “final exam” (subject to Dept. approval.)

* Exam grades will be scaled (if the unscaled average is ≤ 75) to some value in the 75 to 80 range. MAKEUPS WILL BE GIVEN VERY RELUCTANTLY

Homework: Homework problems will be assigned each Tuesday in class and will be due the next Monday. A serious effort should be made on these exercises. It will probably be difficult to completely learn the course material without working problems; also exam questions may bear some resemblance to homework exercises. Additionally, there will be a grade associated with the homework; this will be determined by grading some of the homework. (There is no homework grader available during the Summer, so it won’t be possible to grade all the homework in detail.)

Quizes: Three quizes will be given in the discussion classes on 6/19 (Mon), 7/14 (Fri), and 8/4 (Fri). These will probably be short (about 1/2 hour) and consist of one problem. All three quiz grades will be combined to form one grade counted equal to a single test grade.

Final Grades: In computing a final grade, I will count the three tests, one total homework grade and one total quiz grade equally. (So there will be five grades to average.)